OpenStudy (anonymous):

HELP I WILL MEDAL AND FAN Write the expression as a fraction with a positive exponent. W^-8

OpenStudy (anonymous):

@Simple10 @563blackghost @geny55

OpenStudy (anonymous):

Mortal why do you summon me!

OpenStudy (anonymous):

because this mortal needs help!

OpenStudy (anonymous):

I have never done this math but here is my work. W^-8 W must equal a positive number above 8 to become a positive. W^-8 Plus this equation looks like it is written incorrectly because there is no ^ or there is no base number before ^.

OpenStudy (here_to_help15):

\[\frac{ 1 }{ w ^{8} } \]

OpenStudy (anonymous):

its W to the power of -8 i didnt know how else to write it right when typing @Simple10

OpenStudy (anonymous):

Oh makes sence yea, Like @Here_to_Help15 wrote but anything above 8 would work...

OpenStudy (here_to_help15):

\[\frac{ 1w ^{-8} }{ 1 } = \frac{ 1 }{ w ^{8}}\]

OpenStudy (anonymous):

Sheezus I am only in 9th grade, I only understand the last half of this work XD!

OpenStudy (here_to_help15):

No It is not possible to have a - exponent so we turn into a fraction to make it positive.

OpenStudy (anonymous):

I'm so confused .-. bu thanks @Here_to_Help15